OpenStudy (anonymous):

ok but i must compare the two hamburgers by comparing the volume of the beef in each burger. the volume V of a burger is V= Ah, whee A is the area and h is the height. the area of the lite burger is nr^2, where r is the radius (half the distance across the middle) of the burger. for the lite burger, r= 2 inches, so its area is

OpenStudy (anonymous):

V = Ah \[A = \pi*r^2 = 4*\pi\]